A 40.0 g sample of ethanol releases 2952 J as it cools from 50.0 °C. Calculate the final temperature of the ethanol.


1
Expert's answer
2022-05-06T04:14:03-0400

The link between the energy release and the temperature is:


"Q = c \times m \times \Delta T = c \times m \times (T2-T1)"

where "T_2" and "T_1" are final and initial temperature respectively and "c" is a specific heat capacity (2.46 for ethanol).


The change in temperature is:

"\Delta T = \frac{Q}{c \times m} = \frac{2952}{2.46\times40} = 30 ^oC"


"T2 = \Delta T + T1 = 30+50 = 80 ^oC"